if your scale factor is 1 1/2 inches= 10 ft how much feet would you have if you had 2 1/2 inches how much feet do you have

Answer by checkley79(3341)

1.5/2.5=10/X
1.5X=2.5*10
1.5X=25
X=25/1.5
X=16.67 FEET ANS.